# MIPS at SemEval-2024 Task 3: Multimodal Emotion-Cause Pair Extraction in Conversations with Multimodal Language Models

### v1

@@ 1 Introduction
Emotion recognition in conversations has garnered significant attention due to its wide-ranging applications, such as optimizing customer service interactions and tailoring content recommendations based on user emotions (Yingjian et al., 2023). However, a fundamental challenge lies in identifying the causal determinants of emotional states (Li et al., 2022b). Recent research emphasizes the exploration of causes triggering emotions from multimodal data (Li et al., 2022a), followed by further generation tasks based on multimodal emotional cues (Xu et al., 2024).
@@
Practical conversations often exhibit multimodal cues through facial expressions, vocal changes, and textual content. Recognizing the significance of multimodal information, Wang et al. (2023) proposed the task of Multimodal Emotion-Cause Pair Extraction in Conversations (MECPE) as a critical step towards understanding the fundamental elicitors of emotions. The Emotion-Cause-in-Friends (ECF) dataset, introduced in SemEval 2024 Task 3, incorporates additional modalities such as images and audio alongside the original textual data, enabling a more realistic and comprehensive approach to emotion understanding. Figure 1 illustrates an example of an annotated conversation in the ECF dataset, where variations in facial expressions directly mirror the emotions expressed by the characters.

### v2

@@ 1 Introduction
Emotion-Cause pair extraction in conversations has garnered significant attention due to its wide-ranging applications, such as optimizing customer service interactions and tailoring content recommendations based on user emotions (Xia and Ding, 2019). However, a fundamental challenge lies in identifying the causal determinants of emotional states (Li et al., 2022b). Recent research emphasizes the exploration of causes triggering emotions from multimodal data (Li et al., 2022a), followed by further generation tasks based on multimodal emotional cues (Xu et al., 2024).
@@
Practical conversations often exhibit multimodal cues through facial expressions, vocal changes, and textual content. Recognizing the significance of multimodal information, Wang et al. (2023) proposed the task of Multimodal Emotion-Cause Pair Extraction in Conversations (MECPE) as a critical step towards understanding the fundamental elicitors of emotions. The Emotion-Cause-in-Friends (ECF) dataset, introduced in SemEval 2024 Task 3 (Wang et al., 2024), incorporates additional modalities such as images and audio alongside the original textual data, enabling a more realistic and comprehensive approach to emotion understanding. Figure 1 illustrates an example of an annotated conversation in the ECF dataset, where variations in facial expressions directly mirror the emotions expressed by the characters.